1. What is the volume in L of 4.00 moles of gas produced in a reaction at STP?

            A 22.4             B 5.60             C 89.60           D 0.179                    

2. If the pressure exerted by 50.0mL of gas at 735 mm Hg increases to 925 mm Hg,

       what is the new volume?

A 39.7mL B. 48.4 mL      C.50.0 mL       D. 62.9mL
Answer #1

1. What is the volume in L of 4.00 moles of gas produced in a reaction at STP?

            A 22.4             B 5.60             C 89.60           D 0.179                    

Answer: We know that,

At STP, 1 mole of gas has the volume 22.4 L.

So, 4 moles of gas produced in a reaction will have the volume (22.4)*4 L = 89.60 L

So, the answer is option C.

2. If the pressure exerted by 50.0mL of gas at 735 mm Hg increases to 925 mm Hg,

       what is the new volume?

A 39.7mL B. 48.4 mL      C.50.0 mL       D. 62.9mL

Answer:

Given, intial volume of gas, V1 = 50 mL, inital pressure, P1=735 mm Hg

final pressure, P2=925 mm Hg

Lets assume final volume will be V2.

From the Boyle's law we know,

P1V1=P2V2

So, V2 = (P1*V1)/P2

= (735*50/925) mL

= 39.7 mL

So, option A is correct.